# Effect of the Large Magellanic Cloud on the kinematics of Milky Way satellites and virial mass estimate

### v1

@@ 3.2 Robust halo mass estimator
The mass of a system in equilibrium should be proportional to a measure of the size of the system and the velocity variance of some tracers of the halo potential. We therefore searched for an estimator of this kind, taking into account that during the LMC passage, the distance distribution of the satellites shifts to smaller distances, while velocity distribution shifts to larger velocities. We find that an accurate estimator of the halo mass enclosing the density contrast of 200 relative to the critical mass, $M_{\rm 200c}$, is given by
@@
where the constant $c=2.85\pm 0.028\times 10^{5}\,\,M_{\odot}$ estimated using the true values of $M_{\rm 200c}$ of the 32 Caterpillar host halos, $r_{\rm med}$ is the median halocentric distance of the satellites within $r_{\rm max}=300$ kpc in kiloparsecs, and $\sigma^{2}_{\rm 3D}$ is the variance of $v_{\rm 3D}\times\mathbf{v}_{r}/|v_{r}|$ for satellites within $r_{\rm max}$ in units of kilometers per second. Note that the variance of velocities with the sign of radial velocities is used here, not just the variance of the absolute values of $v_{\rm 3D}$. This variance measures the width of the 3D velocity distribution of halo satellites shown in Figure 5.
@@ 4 Summary and conclusions
    In agreement with results of Erkal et al. (2020), we find that the effect of the passage of an LMC-like satellite on the velocity distribution is substantial and this effect can bias mass estimates based on velocities. We show that during the pericentric passage, the velocity distribution develops a high-$v$ tail due to the reflex motion of the inner part of the halo, while the satellite spatial distribution becomes more centrally concentrated.
@@
    We apply the estimator to the MW satellite system and estimate the virial mass of the Milky Way to be $M_{\rm 200c}=9.96\pm 1.45\times 10^{11}\,\,M_{\odot}$, in good agreement with several recent estimates using other methods.
@@ Acknowledgements
We are grateful to Vasily Belokurov and the UChicago structure formation group for useful discussions during this project. We thank Alexander Ji and the Caterpillar collaboration for providing halo tracks of the Caterpillar simulations used in this study. AK was supported by the National Science Foundation grants AST-1714658 and AST-1911111 and NASA ATP grant 80NSSC20K0512. SW was supported by the University of Chicago CCRF’s Quad Research Scholarship program. Analyses presented in this paper were greatly aided by the following free software packages: NumPy (Van Der Walt et al., 2011), SciPy (Jones et al., 01 ), Matplotlib (Hunter, 2007), and GitHub. We have also used the Astrophysics Data Service (ADS) and arXiv preprint repository extensively during this project and the writing of the paper.

### v2

@@ 3.2 Robust halo mass estimator
The mass of a system in equilibrium should be proportional to a measure of the size of the system and the velocity variance of some tracers of the halo potential. We therefore searched for an estimator of this kind, taking into account that during the LMC passage, the distance distribution of the satellites shifts to smaller distances, while velocity distribution shifts to larger velocities. We find that an accurate estimator of the halo mass enclosing the density contrast of 200 relative to the critical mass, $M_{\rm 200c}$, is given by [^2] Alternatively, one can recast this expression as $M_{\rm 200c,est}=c^{\prime}\,G^{-1}r_{\rm med}\,\sigma^{2}_{\rm 3D}$, where $c^{\prime}=1.226\pm 0.012$ and the gravitational constant $G$, $r_{\rm med}$ and $\sigma^{2}_{\rm 3D}$ are in cgs units.
@@
where the coefficient of proportionality $c=2.85\pm 0.028\times 10^{5}\,\,M_{\odot}$ was estimated using the true values of $M_{\rm 200c}$ of the 32 Caterpillar host halos, $r_{\rm med}$ is the median halocentric distance of the satellites within $r_{\rm max}=300$ kpc in kiloparsecs, and $\sigma^{2}_{\rm 3D}$ is the variance of $v_{\rm 3D}\times\mathbf{v}_{r}/|v_{r}|$ for satellites (i.e. velocities plotted in Fig. 5) within $r_{\rm max}$ in kilometers per second. Note that the variance of velocities with the sign of radial velocities is used here, not just the variance of the absolute values of $v_{\rm 3D}$. This variance measures the width of the 3D velocity distribution of halo satellites shown in Figure 5.
@@ 4 Summary and conclusions
    In agreement with results of Erkal et al. (2020) and Deason et al. (2021), we find that the effect of the passage of an LMC-like satellite on the velocity distribution is substantial and this effect can bias mass estimates based on velocities. We show that during the pericentric passage, the velocity distribution develops a high-$v$ tail due to the motions induced by LMC, while the satellite spatial distribution becomes more centrally concentrated.
@@
    We apply the estimator to the MW satellite system and estimate the virial mass of the Milky Way to be $M_{\rm 200c}=9.96\pm 1.45\times 10^{11}\,\,M_{\odot}$, in good agreement with several recent estimates using other methods (e.g., Erkal et al., 2020; Deason et al., 2021; Vasiliev et al., 2021; Correa Magnus & Vasiliev, 2022).
@@ Acknowledgements
We are grateful to Vasily Belokurov, Denis Erkal and the UChicago structure formation and Cambridge Streams group for useful discussions during this project. We thank Alexander Ji and the Caterpillar collaboration for providing halo tracks of the Caterpillar simulations used in this study. AK was supported by the National Science Foundation grant AST-1911111 and NASA ATP grant 80NSSC20K0512. SW was supported by the University of Chicago CCRF’s Quad Research Scholarship program. Analyses presented in this paper were greatly aided by the following free software packages: NumPy (Van Der Walt et al., 2011), SciPy (Jones et al., 01 ), Matplotlib (Hunter, 2007), and GitHub. We have also used the Astrophysics Data Service (ADS) and arXiv preprint repository extensively during this project and the writing of the paper.